OpenStudy (anonymous):

solve the inequality x^2<2

OpenStudy (nottim):

What do you do to isolate x?

OpenStudy (anonymous):

x?

OpenStudy (lilg132):

i'll give you a clue \[\sqrt{}\]

OpenStudy (nottim):

Big clue there. HOLMEs, you may be on tho something.

OpenStudy (nottim):

ER? You still alive?

OpenStudy (anonymous):

i dont know is it -sqrt2 < x < sqrt2

OpenStudy (nottim):

Make no sense. USe equation, or draw.

OpenStudy (lilg132):

ok let me explain one thing if you have x^2 and you square root it you get rid of the ^2 but when you square root one side you must remember to do it to the other side dont forget to click good answer ! :D

OpenStudy (anonymous):

\[-\sqrt{2}<x<\sqrt{2}\]

OpenStudy (nottim):

Yes